Evaluate the indefinite integral

((1 + t/10)^3) dt

using substitution.

Step 1). u=g(t)=

Step 2). du=

Step 3-5). ∫((1+ t / 10)^3) dt=
Answer #1

Solutions Given, I = S((1+ + 3) et u = g(t) = 1+ 7 Let, Let (2) du = 1 dt at 100 = 1 / dt dt = lodu. Now, putting above subst